# Incorporating Coupling Knowledge into Echo State Networks for Learning Spatiotemporally Chaotic Dynamics

### v1

@@ IV.3 Experimental setup
Given prior knowledge of the coupling structure in the observed benchmark DS, our goal was to reconstruct the underlying dynamics of the simulated time-series data using four ESN models: PGC-ESN, Std-ESN, RandC-ESN, and Paral-ESN. To evaluate the performance of these ESN models across different reservoir sizes, where the size for Paral-ESN represents the total sum of all reservoirs, we adopted an experimental setup as follows. The observed time series has a length of $T=2\cdot 10^{5}$, which is evenly split into training and test datasets. Within the training set, $40\%$ is reserved as a validation set during hyperparameter tuning of the input scaling $\sigma$ and spectral radius $\rho$, using grid search with $D_{\mathrm{temp}}$ as the selection metric. We standardize the observed sequence before feeding it into the ESN model during the training phase. The computed mean and standard deviation are also used to scale the observed warm-up sequence during the prediction phase. All evaluation metrics, except for geometric distance, are computed on the descaled observed and predicted sequences. In Std-ESN and RandC-ESN, the densities $P_{\mathrm{in}}$ and $P$ of the input weight matrix $\mathbf{W}_{in}$ and reservoir weight matrix $\mathbf{W}$ are set to match those of the PGC-ESN, where both densities are determined by the ratio of coupled dimensions to total dimensions within the observed system. In Paral-ESN, we set $P_{\mathrm{in}}=1$ and $P=1$ to ensure full connectivity within each cluster, aligning with the structure of PGC-ESN. Additionally, the group size $G$ and interaction length $I$ are configured such that each reservoir corresponds to and receives the same number of observed state dimensions as in PGC-ESN, or as close as possible when the coupling structure is asymmetric. We assessed the attractor reconstruction using 5 sequences during tuning and 10 sequences during testing. Each experiment was repeated with 5 randomly initialized models for tuning and 10 for testing, and the final results were averaged. All the other hyperparameters are shown in Table 1.
@@ IV.4.1 Lorenz-96 system
The coupling structure of the Lorenz-96 system resulted in a ratio of coupled dimensions to total dimensions of $0.1$. For Paral-ESN, we set $G=1$ and $I=2$. Fig. 2 presents the reconstruction results for a specific target sequence using a reservoir size of 500 for PGC-ESN, Paral-ESN, RandC-ESN, and Std-ESN (from left to right). In Fig. 2(a), we visualize both the target and predicted attractors on the first two-dimensional subspace, where it is evident that the PGC-ESN achieves the best attractor reconstruction. In Fig. 2(b), we present the target and predicted power spectra, showing that the PGC-ESN most accurately reproduces the spectral structure, reinforcing its superior performance in capturing long-term dynamics. In Fig. 2(c), we display contour plots of the first 700 time steps of the target and predicted sequences for all models, along with the NRSE error evolution across all phase-space dimensions. The evolution of the NRMSE averaged over all dimensions is also shown in the lower-left corner. The contour plots reveal that the PGC-ESN best captures the pattern of the target sequence, while the NRSE plot demonstrates that only the PGC-ESN achieves a clear short-term prediction with a low NRSE error up to approximately 2 Lyapunov times. Moreover, the PGC-ESN maintains the lowest NRMSE throughout, further validating its superior predictive accuracy. The numerical results for the evaluated metrics are summarized in Table 2.
@@
System ESN Model $D_{\mathrm{geom}}$ $D_{\mathrm{temp}}$ VPT-1.0
@@
Fig. 3(a) presents a comparison of performance across reservoir sizes, starting from 250 and then ranging from 500 to 4000 in increments of 500. Infinite $D_{\mathrm{geom}}$ values resulting from numerical calculations are omitted from the plots. We observe that the PGC-ESN begins achieving a low $D_{\mathrm{geom}}$ value around 13 and $D_{\mathrm{temp}}$ value around 0.1 at a reservoir size of 500, whereas the other three models do not reach comparable performance until a size of 2000. Beyond this point, all models stabilize at similarly low metric values for long-term prediction. This suggests that the PGC-ESN can effectively reconstruct the Lorenz-96 system with a relatively small reservoir, while the other models require significantly larger reservoirs to achieve similar accuracy. The subplot for VPT-1.0 further highlights the PGC-ESN’s superior short-term predictive performance. It achieves clear short-term predictions beyond a reservoir size of 500, exhibiting noticeably higher VPT-1.0 values above 3 Lyapunov times. In contrast, Paral-ESN and the other two models only begin demonstrating meaningful short-term predictions at reservoir sizes of 1500 and 2000, respectively, but ultimately plateau at lower values. Overall, the proposed PGC-ESN consistently outperforms the other models in long-term attractor reconstruction when the reservoir size is small and in short-term prediction across the entire range of reservoir sizes considered for the Lorenz-96 system. The clustered structure in RandC-ESN, which lacks physical guidance, performs similarly to Std-ESN, indicating that the prior knowledge of the coupling structure in the target system is the key to PGC-ESN’s superior performance.
Figure 4: Evaluation results for the Lorenz-96 system under varied experimental settings. (a) Variation over noise level $\kappa$. (b) Variation over target system’s forcing term $F$. (c) Variation over target system’s nonlinearity parameter $\alpha$.
@@ IV.4.2 Kuramoto–Sivashinsky equation
The numerical results for a specific target sequence using a reservoir size of 2000 for all models are shown in Table 3, while the corresponding reconstruction plots are provided in Appendix. C. These results demonstrate that our proposed PGC-ESN achieves the most accurate long- and short-term predictions for both homogeneous and inhomogeneous KS equations at this reservoir size.
@@
System ESN Model $D_{\mathrm{temp}}$ VPT-1.0
@@
For the homogeneous case shown in Fig. 3 (b), we observe that PGC-ESN begins to achieve stable and low $D_{\mathrm{temp}}$ values around 0.1, whereas Paral-ESN starts to do so at size 3000, and the other two models exhibit unstable values after size 1000. The VPT-1 plot shows that the PGC-ESN achieves the best short-term predictive power across the whole range, saturating at around 4 Lyapunov times after a reservoir size of 3000. In contrast, Paral-ESN and the other two models stabilize at around 3 and 1 Lyapunov times, respectively, after a reservoir size of 5000. A similar trend for the inhomogeneous KS equation is shown in Fig. 3(c). Overall, the proposed PGC-ESN consistently outperforms the other models in both long-term attractor reconstruction and short-term prediction for both the homogeneous and inhomogeneous KS equations across the entire range of reservoir sizes considered.
Figure 5: Investigation of incorporating an imperfect coupling structure. (a) A schematic plot of the PartPGC-ESN with rewired couplings. (b) Evaluation results for the Lorenz-96 system.
@@ IV.5 Evaluation on Lorenz-96 system with varied experimental settings
Based on the evaluation results for learning the Lorenz-96 system across different reservoir sizes, all models reach saturation at a size of 3000. Therefore, we then fixed the reservoir size at 3000 and evaluated the ESN models on the Lorenz96 system under varied experimental settings. The results using the metric $D_{\mathrm{temp}}$ are presented here, while additional metrics can be found in Appendix. C.
First, we analyzed the ESN models’ robustness against noise in the training data by varying the noise levels $\kappa$ while assessing models tuned under the standard setting. Specifically, we vary the noise levels from 00 to $200\text{\textperthousand}$ in increments of $20\text{\textperthousand}$. Fig. 4 (a) presents the evaluation results. We observe that all models’ performance deteriorates as the noise level increases, especially beyond 120. Among them, Paral-ESN exhibits the most gradual decline, followed by PGC-ESN, while the other two models deteriorate more noticeably. This suggests that the coupling structure incorporated in both PGC-ESN and Paral-ESN enhances their robustness against noise. Additionally, in Appendix. A, we present the evaluation results for varying training data sizes.
@@
Next, we investigated the impact of systematic nonlinearity by varying the forcing term $F$ and the nonlinearity parameter $\alpha$ in the Lorenz-96 system, tuning and evaluating all models accordingly. Specifically, we vary the forcing term $F$ from 8 to 44 in increments of 4 and the nonlinearity parameter $\alpha$ from 1 to 10 in increments of 1. As both $F$ and $\alpha$ increase, the system’s maximum Lyapunov exponent (MLE) rises consistently, reaching approximately 9.05 and 14, respectively. Fig. 4 (b) presents the evaluation results when the value of $F$ is varied. The results indicate that the PGC-ESN consistently achieves effective attractor reconstruction across the entire range of $F$. In contrast, the other models begin to deteriorate beyond $F=24$. Fig. 4 (c) presents the evaluation results for variation in $\alpha$, also demonstrating that the PGC-ESN consistently outperforms the other models across the entire range of $\alpha$. Notably, in both cases, the Paral-ESN exhibits a somewhat more moderate deterioration compared to the Std-ESN and RandC-ESN, which do not incorporate prior coupling knowledge.
@@
In conclusion, these results further highlight the advantages of incorporating coupling knowledge into ESN models, enhancing their robustness to noise and enabling superior learning performance as the system’s nonlinearity increases.
@@ IV.6 Evaluation on Lorenz-96 system with imperfect prior coupling knowledge
In real-world applications, prior knowledge of the target system’s coupling structure is often imperfect. To investigate the impact of such imperfections, we introduced perturbations to the reservoir’s coupling structure of our proposed PGC-ESN model. Specifically, we randomly rewired truly coupled clusters and dimensions with a given probability in both the reservoir and input layers for each cluster. We refer to this perturbed model as the Partially Physics-Guided Clustered ESN (PartPGC-ESN). Fig. 5(a) provides a schematic representation of this model, where the yellow arrows indicate the randomly rewired couplings for the $i^{\mathrm{th}}$ cluster, highlighted in green.
@@
We compared the performance of PartPGC-ESN under perturbation probabilities of 0.1, 0.3, and 0.5 with that of the unperturbed PGC-ESN and Std-ESN on the Lorenz-96 system across various reservoir sizes, starting from 250 and then ranging from 500 to 4000 in increments of 500. Fig. 5(b) presents the plots for $D_{\mathrm{temp}}$ and VPT-0.1, while the results for $D_{\mathrm{geom}}$ are provided in Appendix. C. The results indicate that introducing coupling perturbations degrades the performance of PartPGC-ESN compared to that of the unperturbed PGC-ESN. However, with a perturbation probability of 0.1, PartPGC-ESN still outperforms Std-ESN in both long- and short-term predictions, particularly for reservoir sizes between 500 and 2000. As the perturbation probability increases, the performance of PartPGC-ESN further deteriorates, eventually falling behind that of the Std-ESN. These findings suggest that even with imperfect prior coupling knowledge, our proposed approach can still achieve superior long-term attractor reconstruction and short-term prediction compared to Std-ESN, provided the rate of incorrect couplings remains low.

### v2

@@ IV.3 Experimental setup
Given prior knowledge of the coupling structure in the observed benchmark DS, our goal was to reconstruct the underlying dynamics of the simulated time-series data using four ESN models: PGC-ESN, Std-ESN, RandC-ESN, and Paral-ESN. To evaluate the performance of these ESN models across different reservoir sizes, where the size for Paral-ESN represents the total sum of all reservoirs, we adopted an experimental setup as follows. The observed time series has a length of $T=2\cdot 10^{5}$, which is evenly split into training and test datasets. Within the training set, $40\%$ is reserved as a validation set during hyperparameter tuning of the input scaling $\beta$ and spectral radius $\rho$, using grid search with $D_{\mathrm{temp}}$ as the selection metric. We add noise to the observed training sequence and standardize it before feeding it into the ESN model during the training phase. The computed mean and standard deviation are also used to scale the observed warm-up sequence during the prediction phase. All evaluation metrics, except for geometric distance, are computed on the descaled observed and predicted sequences. In Std-ESN and RandC-ESN, the densities $p_{\mathrm{in}}$ and $p$ of the input weight matrix $\mathbf{W}_{in}$ and reservoir weight matrix $\mathbf{W}$ are set to match those of the PGC-ESN, where both densities are determined by the ratio of coupled dimensions to total dimensions within the observed system. In Paral-ESN, we set $p_{\mathrm{in}}=1$ and $p=1$ to ensure full connectivity within each reservoir unit, aligning with the structure of PGC-ESN. Additionally, the group size $G$ and interaction length $I$ are configured such that each reservoir corresponds to and receives the same number of observed state dimensions as in PGC-ESN, or as close as possible when the coupling structure is asymmetric. We assessed the attractor reconstruction using 5 sequences during tuning and 10 sequences during testing. Each experiment was repeated with 5 randomly initialized models for tuning and 10 for testing, and the final results were averaged. All the other hyperparameters are shown in Table 1.
@@ IV.4.1 Lorenz-96 system
The coupling structure of the Lorenz-96 system resulted in a ratio of coupled dimensions to total dimensions of $0.1$. For Paral-ESN, we set $G=1$ and $I=2$. Note that compared to PGC-ESN, which uses the correct coupling structure, each reservoir unit in Paral-ESN receives one additional dimension from the input state on its right-hand side, as illustrated in Fig. 1. Fig. 2 presents the reconstruction results for a specific target sequence using a reservoir size of 500 for PGC-ESN, Paral-ESN, RandC-ESN, and Std-ESN (from left to right). In Fig. 2(a), we visualize both the target and predicted attractors on the first three-dimensional subspace, where it is evident that the PGC-ESN achieves the best attractor reconstruction. In Fig. 2(b), we present the target and predicted power spectra, showing that the PGC-ESN most accurately reproduces the spectral structure, reinforcing its superior performance in capturing long-term dynamics. In Fig. 2(c), we display contour plots of the first 700 time steps of the target and predicted sequences for all models. Fig. 2(d) additionally shows the NRMSE evolution averaged over all state-space dimensions and the NRSE evolution across all dimensions. Note that the NRMSE plot uses a logarithmic scale for error on the $x$-axis to emphasize the error range relevant to short-term prediction. The gray horizontal line indicates an error level of 1 and intersects each model’s error trajectory at its corresponding VPT-1 time. The contour plots of the sequences reveal that the PGC-ESN best captures the pattern of the target sequence, while the NRSE plot demonstrates that only the PGC-ESN achieves a clear short-term prediction with a low NRSE error up to approximately 2 Lyapunov times. Moreover, the NRMSE plot shows that PGC-ESN consistently maintains the lowest error throughout the shown time horizon, with a VPT-1 exceeding 2 Lyapunov times, further validating its superior predictive accuracy. The numerical results for the evaluated metrics, averaged across all sequences and model initializations, are summarized in Table 2.
@@
System ESN Model $D_{\mathrm{geom}}$ $D_{\mathrm{temp}}$ VPT-1
@@
Fig. 3(a) presents a comparison of performance across reservoir sizes, starting from 250 and then ranging from 500 to 4000 in increments of 500. Infinite $D_{\mathrm{geom}}$ values resulting from numerical calculations are omitted from the plots. We observe that the PGC-ESN begins achieving a low $D_{\mathrm{geom}}$ value around 13 and $D_{\mathrm{temp}}$ value around 0.1 at a reservoir size of 500, whereas the other three models do not reach comparable performance until a size of 2000. Beyond this point, all models stabilize at similarly low metric values for long-term prediction. This suggests that the PGC-ESN can effectively reconstruct the Lorenz-96 system with a relatively small reservoir, while the other models require significantly larger reservoirs to achieve similar accuracy. The subplot for VPT-1 further highlights the PGC-ESN’s superior short-term predictive performance. It achieves clear short-term predictions for reservoir sizes greater than 500, exhibiting noticeably higher VPT-1 values above 3 Lyapunov times. In contrast, Paral-ESN and the other two models only begin demonstrating meaningful short-term predictions at reservoir sizes of 1500 and 2000, respectively, but ultimately plateau at lower values. Overall, the proposed PGC-ESN consistently outperforms the other models in long-term attractor reconstruction when the reservoir size is small and in short-term prediction across the entire range of reservoir sizes considered for the Lorenz-96 system. RandC-ESN, which lacks the coupling guidance, performs similarly to Std-ESN, indicating that only introducing a clustered structure in the reservoir layer will not enhance ESN’s learning performance, and knowledge of the coupling structure in the target system is the key to PGC-ESN’s superior performance.
Figure 4: Evaluation results for the Lorenz-96 system under varied experimental settings. (a) Variation over noise level $\kappa$. (b) Variation over sequence temporal resolution. (c) Variation over the target system’s forcing term $F$. (d) Variation over the target system’s nonlinearity parameter $\alpha$.
@@ IV.4.2 Kuramoto–Sivashinsky equation
The numerical results using a reservoir size of 2000 for all models, averaged across sequences and model initializations, are shown in Table 3, while the corresponding reconstruction plots are provided in Appendix D.3. These results demonstrate that our proposed PGC-ESN achieves the most accurate long- and short-term predictions for both homogeneous and inhomogeneous KS equations at this reservoir size.
@@
System ESN Model $D_{\mathrm{temp}}$ VPT-1
@@
For the homogeneous case shown in Fig. 3 (b), we observe that PGC-ESN begins to achieve stable and low $D_{\mathrm{temp}}$ values around 0.1 from size 1000, whereas Paral-ESN starts to do so at size 3000, and the other two models exhibit unstable values until size 5000. The VPT-1 plot shows that the PGC-ESN achieves the best short-term predictive power across the whole range, exceeding 4 Lyapunov times for reservoir sizes greater than 3000. In contrast, Paral-ESN and the other two models exceed 3 and 1 Lyapunov times, respectively, for reservoir sizes greater than 4000. A similar trend for the inhomogeneous KS equation is shown in Fig. 3(c). Overall, the proposed PGC-ESN consistently outperforms the other models in long-term attractor reconstruction when the reservoir size is small and in short-term prediction across the entire range of reservoir sizes considered for both the homogeneous and inhomogeneous KS equations.
Figure 5: Investigation of incorporating imperfect prior coupling knowledge. (a) A schematic plot of the PartPGC-ESN with rewired couplings. (b) Evaluation results for the Lorenz-96 system.
@@ IV.5 Evaluation on Lorenz-96 system with varied experimental settings
Based on the evaluation results for learning the Lorenz-96 system across different reservoir sizes, all models reach saturation at a size of 3000. Therefore, we then fixed the reservoir size at 3000 and evaluated the ESN models on the Lorenz96 system under varied experimental settings. The results using the metric $D_{\mathrm{temp}}$ are presented here, while additional metrics can be found in Appendix D.1.
First, we analyzed the robustness of ESN models to noise in training data and to variations in sequence temporal resolution while assessing models tuned under the standard setting. Specifically, we vary the noise levels $\kappa$ from 0 to $200\text{\textperthousand}$ in increments of $20\text{\textperthousand}$, and simulate varied temporal resolution by downsampling the observed sequences to step sizes ranging from $\Delta t$ to $5\Delta t$ in increments of $\Delta t$, where $\Delta t$ denotes the integration step size used in simulation of the observed trajectory, as described in Subsection IV.1. Fig. 4 (a) presents the evaluation results for varying noise level. We observe that all models’ performance deteriorates as the noise level increases, especially beyond 120‰. Among them, Paral-ESN and PGC-ESN exhibit a more gradual performance decline, while the other two models deteriorate more noticeably. Notably, Paral-ESN achieves the lowest $D_{\mathrm{temp}}$ once the noise level exceeds $\kappa=140\text{\textperthousand}$. We consider that the independent computation of each reservoir’s output layer in Paral-ESN may contribute to its enhanced robustness under high noise conditions. Note that PGC-ESN achieves the best short-term prediction performance across the entire range of noise levels, as shown in Appendix D.1. Overall, the superior robustness of PGC-ESN and Paral-ESN—both of which incorporate prior knowledge of the target system’s coupling structure to some extent—compared to Std-ESN and RandC-ESN, which lack this guidance, suggests that leveraging such physical information plays a key role in mitigating the impact of noise.
Fig. 4(b) shows the evaluation results under varying sequence temporal resolutions. We observe that PGC-ESN consistently achieves the best performance across all step sizes, while Paral-ESN exhibits a more gradual performance decline compared to Std-ESN and RandC-ESN, and outperforms both when the step size reaches $3\Delta t$. Additionally, in Appendix D.1, we present the evaluation results for varying training data sizes.
@@
Next, we investigated the impact of systematic nonlinearity by varying the forcing term $F$ and the nonlinearity parameter $\alpha$ in the Lorenz-96 system, tuning and evaluating all models accordingly. Specifically, we vary the forcing term $F$ from 8 to 44 in increments of 4 and the nonlinearity parameter $\alpha$ from 1 to 10 in increments of 1. As both $F$ and $\alpha$ increase, the system’s maximum Lyapunov exponent (MLE) rises consistently, reaching approximately 9.05 and 14, respectively. Fig. 4 (c) presents the evaluation results when the value of $F$ is varied. The results indicate that the PGC-ESN consistently achieves effective attractor reconstruction across the entire range of $F$. In contrast, the other models begin to deteriorate beyond $F=24$. Fig. 4 (d) presents the evaluation results for variation in $\alpha$, also demonstrating that the PGC-ESN consistently outperforms the other models across the entire range of $\alpha$. Notably, in both cases, the Paral-ESN exhibits a somewhat more moderate deterioration compared to the Std-ESN and RandC-ESN, which do not incorporate prior coupling knowledge.
@@
In conclusion, these results further highlight the advantages of incorporating coupling knowledge of the target system into ESN models—such as in our PGC-ESN and Paral-ESN—enhancing their robustness to varying training noise level, training data size, and sequence temporal resolution, while also enabling superior learning performance as the system’s nonlinearity increases.
@@ IV.6 Evaluation on Lorenz-96 system with imperfect prior coupling knowledge
In real-world applications, prior knowledge of the target system’s coupling structure is often imperfect. To investigate the impact of such imperfections, we introduced perturbations to the reservoir’s coupling structure of our proposed PGC-ESN model. Specifically, when establishing connections to coupled clusters in the reservoir layer or to coupled input units in the input layer for each reservoir cluster, we randomly rewire each connection to a non-coupled and previously unconnected place, with a given perturbation probability. We refer to this perturbed model as the Partially Physics-Guided Clustered ESN (PartPGC-ESN). Fig. 5(a) provides a schematic representation of this model, where the yellow arrows indicate the randomly rewired couplings for the $i^{\mathrm{th}}$ cluster highlighted in green.
@@
We compared the performance of PartPGC-ESN under perturbation probabilities of 0.1, 0.3, and 0.5 with that of the unperturbed PGC-ESN, Std-ESN, and Paral-ESN on the Lorenz-96 system, using the standard settings described in Subsection IV.3. The evaluation is conducted across various reservoir sizes, starting from 250 and then ranging from 500 to 4000 in increments of 500. Fig. 5(b) presents the plots for $D_{\mathrm{temp}}$ and VPT-1, while the results for $D_{\mathrm{geom}}$ are provided in Appendix D.2. The results indicate that introducing coupling perturbations degrades the performance of PartPGC-ESN compared to that of the unperturbed PGC-ESN. However, with a perturbation probability of 0.1, PartPGC-ESN still outperforms Std-ESN and Paral-ESN in long-term attractor reconstruction for reservoir sizes smaller than 2000, and in short-term prediction across the entire range of considered reservoir sizes—despite exhibiting slight instability in a few cases. As the perturbation probability increases, the performance of PartPGC-ESN further deteriorates, eventually falling behind that of the Std-ESN and Paral-ESN. These findings suggest that even with imperfect prior coupling knowledge, our proposed approach can still enhance ESN’s learning performance in long-term attractor reconstruction under limited model capacity, and in short-term prediction even when the model size is sufficiently large, provided the rate of incorrect couplings remains low.
Figure 6: Investigation of incorporating extracted coupling knowledge. (a) Color map of the coupling strength matrices for the Lorenz-96 system inferred using transfer entropy. (b) Corresponding inferred coupling matrices when using reservoir layer density $p=0.1$. (c) Evaluation results for the Lorenz-96 system.
IV.7 Evaluation on Lorenz-96 system without prior knowledge
To enhance the practical applicability of our proposed approach, we now consider a scenario in which no prior knowledge of the target system is available. In this setting, the guidance for designing the coupling structure of AutoPGC-ESN’s reservoir is extracted directly from the training data using transfer entropy. We compare the performance of AutoPGC-ESN with that of Std-ESN on the Lorenz-96 system. In addition to tuning the spectral radius $\rho$ and input scaling $\beta$ over the same range as in the standard setting, we also tune the reservoir layer density $p$ over the set $\{0.05,\ 0.1,\ 0.2,\ 0.3\}$ using grid search, and set the input layer density $p_{\mathrm{in}}$ equal to $p$. This range emphasizes sparse configurations, as sparsity is known to improve the learning performance of ESNs generally. Lukoševičius (2012) The other settings follow the standard setup described in Subsection IV.3. As the ratio of inferred coupled dimensions (including self-coupling) to total dimensions of the target system is set to $p$, the number of inferred coupled dimensions is then automatically determined by the tuning result. The evaluation is conducted across various reservoir sizes, starting from 250 and then ranging from 500 to 4000 in increments of 500. PGC-ESN and Paral-ESN are excluded from this experiment, as they rely on prior knowledge of the target system. For comparison, we also present their performance evaluated under the standard setting, as described in Subsection IV.4.1.
Fig. 6(a) presents heatmaps of the coupling strength matrices inferred for the Lorenz-96 system using transfer entropy, computed on a sequence of length $3\times 10^{3}$ and $10^{5}$ in the training dataset after applying noise addition with noise level $\kappa=5\text{\textperthousand}$ and standardization. In each heatmap, the $i^{\mathrm{th}}$ row represents the inferred coupling strengths from all sender dimensions to the $i^{\mathrm{th}}$ receiver dimension, with self-coupling (i.e., the $i^{\mathrm{th}}$ column) set to zero by definition. When using the sequence of length $3\times 10^{3}$, each receiver dimension receives the strongest inferred coupling from its immediate neighbors, while the coupling strengths from non-adjacent dimensions are weaker and exhibit slight variation in magnitude. In contrast, when using the sequence of length $10^{5}$, the heatmap shows a clearer structure: each receiver dimension receives the dominant coupling strengths from the true coupled dimensions, followed by weaker but similarly valued coupling strengths from their neighboring dimensions. Coupling strengths from all other dimensions are weak and relatively uniform in magnitude.
Fig. 6(b) shows the inferred coupling matrices, where for each receiver dimension (highlighted in green), the inferred coupled dimensions are indicated in blue. The results are obtained using a reservoir layer density of $p=0.1$, which yields three (non-self) inferred coupled dimensions per row, matching the true number of non-self couplings of the target system. Using time series of lengths $3\times 10^{3}$ and $10^{5}$ for coupling extraction, the overall accuracies of correctly identified non-self couplings relative to the ground-truth coupling structure reach $89.17\%$ and $100\%$, respectively. These results indicate that longer training sequences facilitate more accurate inference of the underlying coupling structure, capturing a richer representation of the target system’s intrinsic dependencies.
Fig. 6 (c) presents the evaluation results. AutoPGC-ESN is evaluated using coupling knowledge extracted from training sequences of length $3\times 10^{3}$ and $10^{5}$. When the reservoir size is as small as 250, AutoPGC-ESN, leveraging a longer training sequence for coupling extraction, achieves performance comparable to PGC-ESN and outperforms the variant using a shorter sequence in both $D_{\mathrm{temp}}$ and VPT-1. Using either near-perfect coupling knowledge extracted from sequences of length $10^{5}$ or even imperfect coupling knowledge extracted from a sequence of length $3\times 10^{3}$, AutoPGC-ESN consistently outperforms Std-ESN and Paral-ESN in long-term attractor reconstruction for reservoir sizes below 1500, as well as in short-term prediction across the entire range of reservoir sizes considered, despite exhibiting minor instability in a few instances. These results indicate that the coupling knowledge automatically extracted and utilized by our proposed approach, whether near-perfect or slightly imperfect, can still significantly enhance the ESN’s learning performance in long-term attractor reconstruction under limited model capacity, and in short-term prediction even when the model size is sufficiently large. Overall, these findings highlight the practical utility of our method even when prior knowledge of the target system is unavailable.